OpenStudy (anonymous):

How many different committees can be formed from 6 teachers and 33 students if the committee consists of 3 teachers and 3 students?

OpenStudy (anonymous):

I know the combination formula, but I don't know how to combine 6C3 and 33C3 to get the answer.

OpenStudy (anonymous):

\[6C3=\frac{ 6*5*4 }{ 3*2*1 }=20\] \[33C3=\frac{ 33*32*31 }{ 3*2*1 }=11*16*31=?\]

OpenStudy (anonymous):

you don't "combine" them you multiply them \[\binom{6}{3}\times \binom{33}{3}\]

OpenStudy (anonymous):

you got that?

OpenStudy (anonymous):

Thank you!!!